OpenStudy (anonymous):

f(x) = -4^3 - 15x^2 + 18x - 5 Write equation of the tangent line at x = -1

OpenStudy (anonymous):

@jim_thompson5910 Can you help? c:

jimthompson5910 (jim_thompson5910):

First derive with respect to x f(x) = -4x^3 - 15x^2 + 18x - 5 f ' (x) = -3*4x^2 - 2*15x^1 + 18 f ' (x) = -12x^2 - 30x + 18

jimthompson5910 (jim_thompson5910):

plug in x = -1 into f ' (x) = -12x^2 - 30x + 18 to find the slope of the tangent line at x = -1 this will give you the value of m

jimthompson5910 (jim_thompson5910):

tell me what you get for m

OpenStudy (anonymous):

36

jimthompson5910 (jim_thompson5910):

the slope of the tangent line at x = -1 is m = 36, good

jimthompson5910 (jim_thompson5910):

now tell me what you get when you plug x = -1 into the original function f(x)

OpenStudy (anonymous):

-34

jimthompson5910 (jim_thompson5910):

this means the point (-1, -34) lies on the graph of the function f(x)

jimthompson5910 (jim_thompson5910):

you now have a slope and a point...all the necessary ingredients to find the equation of a line

OpenStudy (anonymous):

y = 36x + 2

jimthompson5910 (jim_thompson5910):

that's correct

jimthompson5910 (jim_thompson5910):

nice work

OpenStudy (anonymous):

Awesome! :O Well, ok. How about this, Find the x value(s) where the slope of the tangent line is 0

jimthompson5910 (jim_thompson5910):

plug in f ' (x) = 0 and solve for x f ' (x) = -12x^2 - 30x + 18 0 = -12x^2 - 30x + 18 -12x^2 - 30x + 18 = 0 ... x = ?? or x = ??

OpenStudy (anonymous):

OHHHHHHHHHHHHHHHHHHHHHHH -3 and -0.5

jimthompson5910 (jim_thompson5910):

one of those is wrong

OpenStudy (anonymous):

\[\frac{ 5\pm \sqrt{5^2 - 4(-2)(3)} }{ 2(-2) }\]

OpenStudy (anonymous):

Is there anything wrong with the quadratic formula?

jimthompson5910 (jim_thompson5910):

you should have (5+7)/(-4) and (5-7)/(-4)

OpenStudy (anonymous):

Oh I got it, the 0.5 isn't negative

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

So if they tell me to find the Max/Min points I just use the x values and plug it back in to the original formula to find y, then I'd have the points.

OpenStudy (jhannybean):

mmhmm

jimthompson5910 (jim_thompson5910):

you have to use the first or second derivative test though

OpenStudy (anonymous):

So max would be (-0.25,0.5) min would be (-3,-86)

OpenStudy (anonymous):

What is that o.o

OpenStudy (jhannybean):

The first derivative tells you where the function is increasing and decreasing using the critical points, or the "zeroes" of a function The second derivative tells you where the function is concave up or down,and the inflection points where the original function changes concavity.

OpenStudy (anonymous):

I'm sorry, but can you dumb it down a bit? Lol

OpenStudy (jhannybean):

lets see...